# Feather the Throttle: Revisiting Visual Token Pruning for Vision-Language Model Acceleration

### v1

@@ 3 Attention-based
$\phi_{\text{KNN}}$ 66% 0.151 0.249 0.260 0.296 0.239 0.606
$\phi_{\text{uniform}}$ 66% 0.206 0.286 0.297 0.333 0.280 0.618
@@
$\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$ 61% 0.291 0.272 0.247 0.277 0.272 0.633
8 Attention-based
$\phi_{\text{original}}$ 56% 0.194 0.235 0.240 0.263 0.233 0.622
$\phi_{\textbf{{-R}}}$ 56% 0.271 0.267 0.264 0.292 0.273 0.635
@@ 8 Attention-based
$\phi_{\text{KNN}}$ 55% 0.154 0.244 0.252 0.294 0.236 0.607
$\phi_{\text{uniform}}$ 55% 0.246 0.310 0.309 0.348 0.303 0.618
@@
$\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$ 50% 0.320 0.359 0.354 0.388 0.356 0.644
Table 1: Evaluating alternative criteria for token pruning after the early LLM layers. For each task and pruning layer, we bold the best result and underline the second-best result. Our main findings include: (1) our proposed RoPE-free criteria $\phi_{\textbf{{-R}}}$ substantially improves pruning performance compared to the original criteria $\phi_{\text{original}}$; (2) pruning later ($K=8$) yields higher performance than pruning earlier ($K=3$); and (3) integrating uniform sampling into the attention-based criteria with $\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$ enhances effectiveness. Using $\textit{R}=0.75$ for $\phi_{\text{original}}$, $\phi_{\textbf{{-R}}}$, and $\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$. See §4.1 for criteria definitions.
To assess the validity of these hypotheses across tasks, we compare the early visual token pruning approach shown to have a non-optimal pruning strategy with a variant that prunes the same non-optimal tokens before they enter the LLM, ensuring no visual information from pruned tokens is retained. As an additional comparison, we also evaluate a text-only model in which all visual tokens are removed before entering the LLM. As shown in Figure 4, we find that across all evaluated tasks, allowing shallow-layer information transfer prior to pruning offers little to no improvement over the setup where tokens are pruned before entering the LLM. This finding suggests that the high performance of early visual token pruning on many tasks does not stem from the effectiveness of visual information transfer in early layers but rather reflects that many benchmarks do not demand a detailed understanding of visual information to answer questions accurately. Note that for the majority of these benchmarks (except VizWiz and AI2D), there is a substantial performance dropoff when not including any visual information. This finding indicates that solely comparing vision enabled and disabled setups does not provide an adequate assessment of how well a benchmark assesses visual grounding.
Finding 3: The majority of evaluated benchmarks do not require fine-grained visual grounding, as they can often be answered using only visual tokens located towards the bottom of the image.
4 Improving Visual Token Pruning
Based on our findings from §3, we now seek to enhance the studied VLM acceleration approach to better preserve fine-grained visual capabilities while still achieving computational efficiency. To this end, we propose alternative pruning criteria and demonstrate that our modifications enable effective pruning even when applied after early LLM layers (§4.1). Guided by our insights, we present our final approach, FEATHER (Fast and Effective Acceleration wiTH Ensemble cRiteria) (§4.2), and showcase its impressive visual capabilities while offering high computational efficiency (§4.3).
4.1 Evaluating pruning criteria
Given that the criteria for token pruning is based on the attention score received from the last text token and attention is dependent on token relative token position distances, the bias towards selecting bottom tokens is unsurprising. Namely, modeling position with RoPE in the LM results in a long-term decay property [27], causing larger attention scores for visual tokens closer to the last text token (refer to the original paper for a formal explanation). As previous work demonstrates, there is greater emphasis on shorter-distance information in more shallow layers[11], explaining why the positional bias is most prevalent when pruning after early layers compared to after later layers.
From this observation, we propose the following simple adjustment to the pruning criteria in order to mitigate this bias and more effectively select tokens relevant to the text instruction. Specifically, we propose the criteria $\phi_{\textbf{{-R}}}$, in which we still compute the attention score received from the last text token, except we do not apply RoPE to the attention mechanism, thereby removing the long-term decay effect. Note that we only remove RoPE for the criteria calculation, as removing positional information from the VLM itself would surely have an adverse effect. In addition to this altered criteria, we also explore the following criteria that are not based on attention.
$\phi_{\text{KNN}}$: Inspired by dynamic visual tokenization [13], in this criteria, we select tokens based on their local density. Specifically, given the visual tokens $z_{img}$ from the vision backbone, we compute local density $\rho_{i}$ for each token $z_{img}[i]$ where $i\in\{0,1,...,n-1\}$ using K-nearest neighbors. We then calculate the distance index $\delta_{i}$ for each token $z_{img}[i]$ as:
@@ 4.1 Evaluating pruning criteria
This distance index represents how far away $z_{img}[i]$ is from other high-density tokens. With this, we use $\rho_{i}*\delta_{i}$ as the token importance score for the criteria. To match the computational savings of $\phi_{\text{uniform}}$, we select 196 tokens.
In Table 1, we report results evaluating the effectiveness of these criteria, particularly focusing on the challenging vision-centric localization task. Comparing $\phi_{\textbf{{-R}}}$ with $\phi_{\text{original}}$, we find that by removing RoPE, at $K=3$ there is a 183% average improvement on localization tasks and at $K=8$, there is a 17% average improvement. These performance improvements demonstrate that once the impact of token position on attention is removed, attention score can more effectively be used for the criteria when applied after early LLM layers. Note that the narrowing gap in performance between $\phi_{\textbf{{-R}}}$ and $\phi_{\text{original}}$ from $K=3$ to $K=8$ is likely because $\phi_{\text{original}}$ has less of a bias towards selecting bottom image tokens as the pruning layer increases.
Insight 1: Once removing the criteria tendency of selecting bottom image tokens, pruning after early LLM layers becomes substantially more effective.
@@
3 0.238 0.162 0.145 0.163
8 0.267 0.298 0.268 0.298
Table 2: Assessing localization performance when using the selected tokens from $\phi_{\textbf{{-R}}}$ for the entirety of the LLM. We find that $\phi_{\textbf{{-R}}}$ applied at a later LLM layer results in a better selection of tokens.
Comparing $\phi_{\textbf{{-R}}}$ when $K=8$ versus $K=3$, we see that the average localization performance improves by 63%. However, it remains unclear whether this performance increase is affected by factors other than the ability of the criteria to select important tokens. Therefore, to directly compare the criteria across different layers, we take the tokens selected by both criteria and evaluate how well the model can perform with only these tokens passed into the LLM (as is done in §3.4). As shown in Table 2, we find the selected tokens from $\phi_{\textbf{{-R}}}$ applied after layer 8 are superior compared the tokens from $\phi_{\textbf{{-R}}}$ applied after layer 3.
Insight 2: Even with the enhanced criteria for pruning after shallow LLM layers, pruning later still improves the criteria and downstream performance.
@@
When comparing the attention-based criteria $\phi_{\textbf{{-R}}}$ with the two non-attention based criteria $\phi_{\text{KNN}}$ and $\phi_{\text{uniform}}$, we find that $\phi_{\textbf{{-R}}}$ results in performance on the OCID-Ref dataset, while $\phi_{\text{KNN}}$ and $\phi_{\text{uniform}}$ outperform on the RefCOCO tasks. One possible explanation for the varying effectiveness of different criteria is that, since OCID-Ref evaluates localization in cluttered environments, incorporating information from the text instruction into the criteria is crucial. In contrast, RefCOCO tasks contain less cluttered scenes, where broad coverage over the whole image may be most beneficial.
Based on this observation that the attention-based criteria $\phi_{\textbf{{-R}}}$ and non-attention-based criteria $\phi_{\text{uniform}}$ have different strengths and weaknesses, we propose the ensemble criteria $\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$. In this criteria, we utilize $\phi_{\textbf{{-R}}}$ with the same ratio of selected tokens and additionally add a small amount uniformly sampled tokens (using a stride of three instead of two). We find that when $K=3$, this approach improves upon $\phi_{\textbf{{-R}}}$ by 63% but has a slight decrease compared to $\phi_{\text{uniform}}$ by 2.9%. However, when $K=8$, this approach far outperforms both of the individual criteria, improving $\phi_{\textbf{{-R}}}$ by 30% and $\phi_{\text{uniform}}$ by 17%. Note that this setup results in slightly less computation efficiency than $\phi_{\textbf{{-R}}}$ (7% drop in FLOPS reduction for $K=3$ and 6% for $K=16$), but this is outweighed by the performance improvements.
Insight 3: Integrating uniform sampling into the attention-based pruning criteria enhances its effectiveness in early layers.
We provide qualitative examples of various criteria token selection in the supplement.
@@ 4.2 Distilling insights
Guided by our insights, we now present our final approach FEATHER (Fast and Effective Acceleration wiTH Ensemble cRiteria). In this approach, we first perform pruning after an early layer ($K=8$), utilizing our proposed criteria $\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$ to retain $(1-R)\%$ of tokens. Given our finding that the criteria improve as the LLM layer increases, we additionally prune a second time at $K=16$. At this stage, we utilize $\phi_{\textbf{{-R}}}$ as the criteria, as the uniform sampling becomes less effective compared to the attention-based strategy as the pruning layer increases. For the reduction ratio, we choose to only retain $(1-R)^{2}\%$ of the remaining tokens since the attention-based criteria has proved highly effective when pruning at later layers, even when using the suboptimal criteria $\phi_{\text{original}}$ (see Figure 3(c)).
@@ 4.3 Comparison against FastV and PyramidDrop
5 Discussion
In this work, we examine the visual capabilities of the VLM acceleration approach of pruning visual tokens after shallow LLM layers. While strong performance is maintained on most evaluated tasks, it fails on the vision-centric task of localization due to its flawed pruning criteria that predominately selects visual tokens from the bottom part of the image. Observing this same behavior on other tasks, we show that strong performance is largely due to the benchmarks’ inability to assess fine-grained visual capabilities. Next, we propose and evaluate several alternative criteria to improve visual capabilities, ultimately arriving at our final method, FEATHER (Fast and Effective Acceleration wiTH Ensemble cRiteria). This approach first prunes after an early layer using a modified version of the criteria that addresses the issue of bottom token selection while also incorporating uniform sampling to ensure coverage across all image regions. Subsequently, it performs more extensive pruning at a later layer, where the pruning criteria become more effective. We find that FEATHER has more than $\mathbf{5\times}$ performance improvement on localization compared to the original acceleration approach.
While we study a particular VLM architecture and acceleration approach, our findings highlight a broader issue with the evaluation of VLMs. Building on [28] which finds that text-only models can perform well on some vision-language benchmarks, we demonstrate that even benchmarks with a substantial difference in vision enabled and disabled setups may not assess fine-grained visual capabilities. In our work, we address this benchmarking issue by utilizing the vision-centric localization task. However, this task only assesses one particular type of skill. To accurately assess a wide range of visual capabilities, future work may explore how to resolve current dataset biases that models can exploit. Future work can also investigate the capabilities of other VLM acceleration approaches to determine whether they have pitfalls undetected by the current evaluation benchmarks.
@@ 5 Discussion
6 Token Pruning Visualizations
@@
In this supplemental material section, we provide a qualitative analysis comparing the pruning effectiveness of various criteria as well as the final approaches of FEATHER, FastV, and PyramidDrop. Namely, we visualize the ability of approaches to retain important tokens, particularly for localization. In Figure 6 and Figure 7, we visualize pruning from the various criteria assessed in the main text when pruning is done after layers three and eight, respectively. In Figure 8, we visualize pruning from the final approaches of FEATHER, FastV, and PyramidDrop.
6.1 Comparing pruning criteria
@@
We first visualize the retained tokens of various criteria when pruning is applied after layer three (see Figure 6) and layer eight (see Figure 7). We see that these visualizations support our quantitative results from the main paper. Specifically, (1) $\phi_{\textbf{{-R}}}$ removes the criteria tendency of selecting bottom image tokens, resulting in an improved selection of maintained tokens; (2) the attention-based criteria improve when pruning after a later layer; and (3) adding uniform sampling to the attention-based pruning criteria with $\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$ improves token selection.
@@
Figure 6: Visualizing the ability of various pruning criteria to maintain visual tokens relevant to the reference expression when applied after layer three. We observe that $\phi_{\textbf{{-R}}}$ resolves $\phi_{\text{original}}$’s tendency of selecting bottom image tokens and that uniform sampling is a robust approach that improves the token selection effectiveness of $\phi_{\textbf{{-R}}}$ with $\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$. See the main text for criteria definitions.
@@
Figure 7: Visualizing the ability of various pruning criteria to maintain visual tokens relevant to the reference expression when applied after layer eight. We observe that the attention-based criteria are more effective when pruning after this layer compared to after layer three. See the main text for criteria definitions.
6.2 Comparing FEATHER to FastV and PyramidDrop
@@
As shown in Figure 8, when comparing the remaining tokens used for prediction (after layer 16 for FEATHER, layer 24 for PyramidDrop, and layer three for FastV), we see that our approach retains substantially more tokens around and inside the reference expression bounding box.
Figure 8: Visualizing the ability of FEATHER, FastV, and PyramidDrop to retain visual tokens relevant to the reference expression. We observe that our approach retains a substantially higher portion of tokens relevant to the reference expression.

### v2

@@ 4.1 Assessing potential of attention-based criteria
$\phi_{\text{KNN}}$ 66% 23.9 15.1 24.9 26.0 29.6 58.4 39.9 60.9 74.4 58.4 62.8 81.2 55.9 61.5 52.8
$\phi_{\text{uniform}}$ 66% 28.0 20.6 28.6 29.7 33.3 59.0 41.4 61.8 75.9 57.1 64.6 85.2 58.1 61.9 53.0
@@
$\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$ (Ours) 61% 27.2 29.1 27.2 24.7 27.7 61.2 46.6 62.3 77.4 58.4 65.4 86.0 58.9 62.7 54.0
Attention-based
$K=8$ $\phi_{\text{original}}$ 56% 23.3 19.4 23.5 24.0 26.3 59.8 45.0 60.3 76.1 57.8 64.6 85.4 57.5 62.6 53.0
$\phi_{\textbf{{-R}}}$ (Ours) 56% 27.3 27.1 26.7 26.4 29.2 61.4 49.0 61.5 77.4 57.8 65.5 86.7 58.6 63.0 53.7
$\Delta$ +4.0 +7.6 +3.2 +2.5 +2.9 +1.6 +4.0 +1.2 +1.3 +0.0 +0.8 +1.3 +1.1 +0.4 +0.6
@@
$\phi_{\text{KNN}}$ 55% 23.6 15.4 24.4 25.2 29.4 58.6 40.2 61.1 74.5 58.5 62.9 81.4 56.2 60.9 53.0
$\phi_{\text{uniform}}$ 55% 30.3 24.6 31.0 30.9 34.8 59.3 42.2 61.8 76.0 57.4 64.4 85.3 57.9 61.0 53.2
@@
$\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$ (Ours) 50% 35.6 32.0 35.9 35.4 38.8 62.7 51.7 62.4 78.1 58.6 66.0 87.4 59.1 63.6 54.0
@@
3 23.8 16.2 14.5 16.3
8 26.7 29.8 26.8 29.8
Table 2: Assessing localization performance when using the selected tokens from $\phi_{\textbf{{-R}}}$ for the entirety of the LLM. We find that $\phi_{\textbf{{-R}}}$ applied at a later LLM layer results in a better token selection.
Token selection improves when pruning later: Comparing $\phi_{\textbf{{-R}}}$ when $K=8$ versus $K=3$, we see that the average localization performance improves by 63%. However, it remains unclear whether this performance increase is affected by factors other than the ability of the criteria to select important tokens. Therefore, to directly compare the criteria across different layers, we take the tokens selected by both criteria and evaluate how well the model can perform with only these tokens passed into the LLM (as is done in §3.4). As shown in Table 2, we find the selected tokens from $\phi_{\textbf{{-R}}}$ applied after layer 8 are superior compared the tokens from $\phi_{\textbf{{-R}}}$ applied after layer 3.
Integrating uniform sampling with attention-based criteria is beneficial in early layers: When comparing the attention-based criteria $\phi_{\textbf{{-R}}}$ with the two non-attention based criteria $\phi_{\text{KNN}}$ and $\phi_{\text{uniform}}$, we find that $\phi_{\textbf{{-R}}}$ results in better performance on some tasks (e.g., OCID-Ref, TextVQA), while $\phi_{\text{KNN}}$ and $\phi_{\text{uniform}}$ outperform on other tasks (e.g., RefCOCO). This varying effectiveness of criteria types presumably comes from some tasks requiring a more detailed understanding of specific image regions while others benefit from a broader understanding with full image coverage. When we combine these two criteria types with $\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$, we find that when $K=3$ this approach improves upon localization performance of $\phi_{\textbf{{-R}}}$ by 63% but has a slight decrease compared to $\phi_{\text{uniform}}$ by 2.9%. However, when $K=8$, this approach far outperforms both of the individual criteria, improving $\phi_{\textbf{{-R}}}$ by 30% and $\phi_{\text{uniform}}$ by 17%. Note that this setup results in slightly less computation efficiency than $\phi_{\textbf{{-R}}}$ (7% drop in FLOPS reduction for $K=3$ and 6% for $K=16$), but this is outweighed by the performance improvements.
We provide qualitative examples of various criteria token selection and additional results using a different model setup in the supplement.
@@ 4.2 Distilling insights
Guided by our insights, we now present our final approach FEATHER (Fast and Effective Acceleration wiTH Ensemble cRiteria). In this approach, we first perform pruning after an early layer ($K=8$), utilizing our proposed criteria $\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$ to retain $(1-R)\%$ of tokens. Given our finding that the criteria improves as the LLM layer increases, we additionally prune a second time at $K=16$. At this stage, we utilize $\phi_{\textbf{{-R}}}$ as the criteria, as uniform sampling should not be necessary when the attention-based criteria reliably selects important tokens. For the reduction ratio, we choose to only retain $(1-R)^{2}\%$ of the remaining tokens since the attention-based criteria has proved highly effective when pruning at later layers, even when using the original criteria $\phi_{\text{original}}$ (see Figure 3(c)).
@@ 4.3 Comparison against FastV and PyramidDrop
5 Conclusion
In this work, we examine the visual capabilities of the VLM acceleration approach of pruning visual tokens after shallow LLM layers. While strong performance is maintained on most evaluated tasks, it fails on more vision-centric tasks like TextVQA and localization due to its flawed pruning criteria that predominately selects visual tokens from the bottom part of the image. Observing this same flawed criteria on other tasks, we show that strong performance is largely due to the benchmarks’ inability to assess fine-grained visual capabilities. Next, we propose and evaluate several alternative criteria to improve visual capabilities, ultimately arriving at our final method, FEATHER. This approach refines the attention-based criteria to address the token selection bias while using uniform sampling for better image coverage. It then prunes more aggressively when the criteria more effectively identifies important tokens. We find that FEATHER has more than $\mathbf{5\times}$ performance improvement on localization compared to the original acceleration approach.
While we study a particular type of VLM acceleration, our work highlights a broad challenge in evaluating VLMs. Building on [34], which shows that text-only models can perform well on some vision-language benchmarks, we demonstrate that even benchmarks with substantial differences between vision-enabled and disabled setups may not assess fine-grained visual capabilities. We address this issue by focusing on more vision-centric localization tasks, though this limits the analysis to a few specific types of skills. To accurately assess a wider range of visual capabilities, future work may explore how to resolve current dataset biases that models can exploit [21]. Additionally, while removing RoPE for token pruning effectively eliminates positional bias, it may introduce unintended effects on attention weights influencing token selection. Future research could investigate more robust methods for encoding positional information in visually conditioned language models to prevent positional artifacts in cross-modal interactions.
@@ 5 Conclusion
A1 Results on Different Model Setup
We additionally experiment with using a DINOv2 + SigLIP visual encoder. As shown in Table A1, we observe the same behavior that removing RoPE substantially improves performance and incorporating uniform sampling is strong.
A2 Additional FEATHER Results
We compare FEATHER performance against FastV and PyramidDrop on all evaluated benchmarks in Table A2.
Localization Open-Ended VQA Challenge Sets
Criteria
FLOPS Red X
Avg
OCID-Ref
RefCOCOg
RefCOCO+
RefCOCO
Avg
TextVQA
GQA
VQAv2
VizWiz
Avg
POPE
TallyQA
VSR
AI2D
Attention-based
$\phi_{\text{original}}$ 68% 27.2 21.9 27.7 27.8 31.1 56.6 35.6 59.1 74.0 57.7 66.1 84.6 60.2 67.1 52.7
$\phi_{\textbf{{-R}}}$ 68% 37.2 37.0 38.7 34.9 38.1 60.1 45.4 60.4 76.5 58.2 66.3 85.9 61.1 65.1 52.9
$\Delta$ +10.0 +15.0 +11.0 +7.0 +7.0 +3.5 +9.7 +1.2 +2.5 +0.5 +0.1 +1.3 +0.9 -2.0 +0.2
Non-attention-based
$\phi_{\text{KNN}}$ 66% 20.5 13.4 22.1 22.0 24.6 54.2 29.9 60.0 70.2 57.0 60.5 77.7 51.9 61.9 50.7
$\phi_{\text{uniform}}$ 66% 38.3 32.7 38.8 38.8 42.7 58.3 37.6 61.9 75.8 58.0 65.8 85.9 60.2 65.0 52.2
Ensemble
$\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$ (Ours) 61% 46.3 41.6 47.3 46.0 50.1 61.3 46.8 62.0 77.7 58.7 66.8 86.9 61.6 65.4 53.3
Table A1: Evaluating criteria using DINOv2 + SigLIP visual encoder. For each task, we bold the best result and underline the second-best result. Using $K=3$ for all setups.
Localization Open-Ended VQA Challenge Sets
Method
FLOPS Red X
GPU Hours
Avg
OCID-Ref
RefCOCOg
RefCOCO+
RefCOCO
Avg
TextVQA
GQA
VQAv2
VizWiz
Avg
POPE
TallyQA
VSR
AI2D
Baseline 0% 20.3 53.2 40.7 56.3 55.0 60.9 64.1 54.9 63.3 78.9 59.3 66.1 87.4 59.3 63.3 54.3
FastV 68% 15.1 5.9 5.7 5.1 6.1 6.7 54.8 31.8 58.4 72.7 56.3 64.0 83.2 57.1 63.3 52.4
PyramidDrop 65% 15.7 28.9 24.0 29.2 29.7 32.9 60.8 47.1 61.2 76.9 57.9 65.3 86.6 58.2 63.4 53.1
FEATHER 64% 15.7 39.3 33.1 40.1 39.7 44.1 61.9 51.4 61.8 77.9 56.5 66.1 87.7 59.1 63.4 54.2
FastV 45% 16.8 29.1 17.5 29.5 33.1 36.1 61.0 45.8 62.3 77.4 58.4 65.7 86.8 59.2 63.3 53.5
PyramidDrop 46% 16.8 46.6 37.4 48.3 47.8 53.0 63.7 53.8 63.1 78.7 59.1 66.2 87.5 59.4 63.5 54.3
FEATHER 48% 16.5 49.7 39.3 52.1 50.9 56.7 63.9 54.6 63.2 78.8 59.0 66.3 87.7 59.2 64.0 54.6
Table A2: Comparing FEATHER performance against FastV and PyramidDrop. The best results are bolded (excluding the baseline method).
In addition, we show performance with respect to total runtime on a NVIDIA L40S in Figure A1.
Figure A1: Total runtime on L40S vs. performance for FastV, PyramidDrop, and FEATHER.
A3 Comparison Against FasterVLM and
VisionZip
We present FasterVLM and VisionZip performance in Table A3. We find that these approaches, while performing comparably to our approach on some benchmarks, perform vastly worse on localization benchmarks. We expect this is because positional information is not maintained in these methods, as image tokens are filtered without altering the positional embeddings. We verify the importance of positional embeddings in §A4. Note that since our setup uses the SigLIP encoder, for FasterVLM (which relies on [CLS] attention), we use the proposed solution in VisionZip of averaging attention each token receives from all others in the sequence.
A4 Token Shuffling Ablation
To assess the impact of positional embeddings on model performance, we shuffle positional embeddings for the image tokens and evaluate both the original VLM and our FEATHER approach. As shown in Table A3, the localization performance of both methods drops drastically for localization tasks, substantially for TextVQA, and relatively little for other benchmarks. This result supports our key insight that many vision-language benchmarks inadequately capture the shortcomings of efficiency methods due to their limited ability to assess fine-grained visual capabilities, particularly for visual grounding.
Localization Open-Ended VQA Challenge Sets
Method
FLOPS Red X
Avg
OCID-Ref
RefCOCOg
RefCOCO+
RefCOCO
Avg
TextVQA
GQA
VQAv2
VizWiz
Avg
POPE
TallyQA
VSR
AI2D
Baseline 0% 53.2 40.7 56.3 55.0 60.9 64.1 54.9 63.3 78.9 59.3 66.1 87.4 59.3 63.3 54.3
Baseline (pos shuffled) 0% 8.0 9.0 7.8 7.1 8.0 59.2 44.1 60.3 75.8 56.8 63.3 86.6 55.3 59.2 51.9
FasterVLM 65% 5.7 8.0 5.9 4.2 4.7 60.9 50.9 59.9 76.5 56.4 66.6 85.2 62.6 63.7 54.7
VisionZip 65% 8.5 7.3 9.0 8.1 9.5 61.1 50.8 60.2 76.7 56.7 66.5 85.3 62.9 63.7 54.3
FEATHER 64% 39.3 33.1 40.1 39.7 44.1 61.9 51.4 61.8 77.9 56.5 66.1 87.7 59.1 63.4 54.2
FEATHER (pos shuffled) 64% 5.3 5.3 4.8 5.2 5.8 57.8 41.7 58.9 75.0 55.5 63.2 86.0 55.7 58.8 52.5
Table A3: Comparison against FasterVLM and VisionZip and positional embeddings ablation (where image token positions are shuffled). The best results are bolded.
A5 Token Pruning Visualizations
@@
In this supplemental material section, we provide a qualitative analysis comparing the pruning effectiveness of various criteria as well as the final approaches of FEATHER, FastV, and PyramidDrop. Namely, we visualize the ability of approaches to retain important tokens, particularly for localization. In Figure A2 and Figure A3, we visualize pruning from the various criteria assessed in the main text when pruning is done after layers three and eight, respectively. In Figure A4, we visualize pruning from the final approaches of FEATHER, FastV, and PyramidDrop.
A5.1 Comparing pruning criteria
@@
We first visualize the retained tokens of various criteria when pruning is applied after layer three (see Figure A2) and layer eight (see Figure A3). We see that these visualizations support our quantitative results from the main paper. Specifically, (1) $\phi_{\textbf{{-R}}}$ removes the criteria tendency of selecting bottom image tokens, resulting in an improved selection of maintained tokens; (2) the attention-based criteria improve when pruning after a later layer; and (3) adding uniform sampling to the attention-based pruning criteria with $\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$ improves token selection.
@@
Figure A2: Visualizing the ability of various pruning criteria to maintain visual tokens relevant to the reference expression when applied after layer three. We observe that $\phi_{\textbf{{-R}}}$ resolves $\phi_{\text{original}}$’s tendency of selecting bottom image tokens and that uniform sampling is a robust approach that improves the token selection effectiveness of $\phi_{\textbf{{-R}}}$ with $\phi_{\textbf{{-R}}}+\phi_{\text{uniform}}$. See the main text for criteria definitions.
@@
Figure A3: Visualizing the ability of various pruning criteria to maintain visual tokens relevant to the reference expression when applied after layer eight. We observe that the attention-based criteria are more effective when pruning after this layer compared to after layer three. See the main text for criteria definitions.
A5.2 Comparing FEATHER to FastV and PyramidDrop
@@
As shown in Figure A4, when comparing the remaining tokens used for prediction (after layer 16 for FEATHER, layer 24 for PyramidDrop, and layer three for FastV), we see that our approach retains substantially more tokens around and inside the reference expression bounding box.
Figure A4: Visualizing the ability of FEATHER, FastV, and PyramidDrop to retain visual tokens relevant to the reference expression. We observe that our approach retains a substantially higher portion of tokens relevant to the reference expression.